OpenStudy (anonymous):

a2 + 13a + 40 over a2 − 3a − 40

OpenStudy (saifoo.khan):

Factor numerator & denomiator

OpenStudy (saifoo.khan):

\[\frac{(x+5)(x+8)}{(a-8)(a+5)}\]

OpenStudy (saifoo.khan):

\[\frac{\cancel{(x+5)}(x+8)}{(a-8)\cancel{(a+5)}}\]

OpenStudy (saifoo.khan):

\[\frac{x+8}{a-8}\]